OpenStudy (anonymous):

Triangle ABC with endpoints A(-1,4), B(3,1) and C(0,-3) is a right Using the principles of Slope, please verify this.

jimthompson5910 (jim_thompson5910):

Find the slopes of the following segments AB, BC, and AC Then multiply each pair of slopes. If you get -1 as a result, then these two line segments will be perpendicular (and form a right angle)

jimthompson5910 (jim_thompson5910):

So hypothetically say you get slope of AB = 2 slope of BC = 3 slope of AC = -1/2 notice how (slope of AB)*(slope of AC) = 2*(-1/2) = -1 So this would show that line segments AB and AC are perpendicular and form a right angle.